# Manuel Morante Kristian Frølich Naveed ur Rehman

### v1

@@ Abstract
Human brains exhibit highly organized multiscale neurophysiological dynamics. Understanding those dynamic changes and the neuronal networks involved is critical for understanding how the brain functions in health and disease. Functional Magnetic Resonance Imaging (fMRI) is a prevalent neuroimaging technique for studying these complex interactions. However, analyzing fMRI data poses several challenges. Furthermore, most approaches for analyzing Functional Connectivity (FC) still rely on preprocessing or conventional methods, often built upon oversimplified assumptions. On top of that, those approaches often ignore frequency-related information despite evidence showing that fMRI data contain rich information that spans multiple timescales. This study introduces a novel methodology, Multiscale Functional Connectivity (MFC), to analyze fMRI data by decomposing the fMRI into their intrinsic modes, allowing us to separate the neurophysiological activation patterns at multiple timescales while separating them from other interfering components. Additionally, the proposed approach accounts for the natural nonlinear and nonstationary nature of fMRI and the particularities of each individual in a data-driven way. We evaluated the performance of our proposed methodology using three fMRI experiments. Our results demonstrate that our novel approach effectively separates the fMRI data into different timescales while identifying highly reliable functional connectivity patterns across individuals. In addition, we further extended our knowledge of how the FC for these three experiments spans among different timescales.
keywords:
fMRI , Multivariate Mode Decomposition , Functional Connectivity , Multiscale
journal: ArXiv
@@
[…]
[inst1]organization=Department of Electrical and Computer Engineering, addressline=Aarhus University, postcode=8200, country=Denmark
1 Introduction
Even at rest, the brain exhibits intricate yet highly organized neurophysiological interactions that control and regulate our body’s main cognitive and physiological functions (Honey et al., 2007; Fair et al., 2009; Poldrack et al., 2011). Furthermore, although the constant interaction between different neuronal networks -along with endogenous somatic interactions- dominates the evolution of brain dynamics, empirical evidence has shown that environmental factors also affect this process by modulating or inducing changes in neuronal activity (Fox et al., 2005; Friston, 2011). As a result, brain activity displays highly complex spatiotemporal dynamic (Bolton et al., 2020b; Lurie et al., 2020).
The complexity of brain activity can also be seen in the ability of the brain to integrate and process a vast amount of information at different timescales, as discussed by Preti et al. (2017). Additionally, neuronal communities often exhibit complex interaction regardless of their anatomical proximity (Honey et al., 2007; Xu et al., 2016). All these natural features render complex spatial and temporal dynamics, which lead to the emergence of sophisticated cognitive processes such as perception, attention, or memory (Carpenter et al., 2021) and, ultimately, shape our thoughts, behaviors, and experiences, even the way we perceive and interpret our past (Schacter and Coyle, 1995).
@@ 1 Introduction
Researchers have used a variety of neuroimaging techniques to study brain activity. Among all those techniques, functional Magnetic Resonance Imaging (fMRI) stands as one of the most popular for both neurophysiological, e.g., (Ozcelik and VanRullen, 2023; Morante et al., 2021; Chen et al., 2020; Chatzichristos et al., 2020), and clinical research, e.g., (Peng et al., 2021; Hannanu et al., 2020; Seo et al., 2019). Unlike other alternative noninvasive neuroimaging techniques, such as Electroencephalography (EEG), e.g., (Singh and Krishnan, 2023), or Magnetoencephalography (MEG) (Koshev et al., 2021), fMRI records brain activity indirectly by measuring the variation in the oxygenation levels on small volumes of tissue, referred to as voxels (Power et al., 2011), due to the metabolic oxygen consumption of the neurons, which is a process referred to as Bold Oxygenation Level-Dependent (BOLD) contrast Poldrack et al. (2011). Furthermore, the fMRI data contains a mixture of several interfering signals beyond brain-induced activity, such as movement, respiratory and cardiac pulsations (Bianciardi et al., 2009; Logothetis and Wandell, 2004).
The majority of fMRI studies, especially in the early days of fMRI research, have focused on identifying both the spatial areas associated with brain activity and their corresponding activation patterns, often in response to specific stimuli or tasks, usually referred to as sources (Friston, 2011; Fair et al., 2009). Methods such as the General Linear Model (GLM) (Poldrack et al., 2011), and matrix factorization techniques, such as Independent Component Analysis (ICA) or Dictionary Learning, stand as some of the most relevant conventional techniques (Morante, 2021). Nonetheless, those simple models are limited, as they lack the capability of exploring more complex interactions among neuronal networks and their Functional Connectivity (FC) (Bolton et al., 2020b).
In consequence, during the last two decades, the study of the FC and dynamic interaction of neuronal networks using fMRI has steadily gained more attention (Lurie et al., 2020), as Friston (2011) envisioned back in 2011. Researchers have steadily proposed a wide range of techniques to reveal the intricate nature of the brain’s FC (Bolton et al., 2020b; Lurie et al., 2020). Many of the most recent approaches – greatly inspired by the success of Machine Learning (ML) and Artificial Intelligence (AI) in a wide range of complicated problems – leverage machine or deep learning-based models due to their potential to uncover hidden interactions among brain networks that are not directly observable from the data (Kim et al., 2021).
No doubt, ML and DL can provide novel and powerful tools that appear well-suited for capturing FC and brain behavior. The literature (Bolton et al., 2020b; Pervaiz et al., 2020), however, shows that most of these approaches still use conventional preprocessing steps or rely on classical fMRI analysis techniques, such as ICA, as a critical preliminary step in their application. Therefore, even if all those novel cutting-edge techniques have great potential, their application will be hindered by the inherent assumptions and limitations of these preliminary steps.
@@
For instance, conventional motion correction approaches rely on the parametric rigid-body model, which assumes that the brain only undergoes rigid motion. These approaches provide a reasonable estimation for main motion effects. Nonetheless, these motion components are richer than this simple model, as they can carry more information regarding certain behavioral and physiological aspects of the participants, such as arousal(Gu et al., 2019), and even age, as shown by Bolton et al. (2020a). Furthermore, these simple models overlook other more subtle sources of motion, such as respiratory or cardiac pulsations. These additional motion components are often challenging to identify since they appear highly structured, individual dependent, and even mimic other brain sources, as evidenced by Chen et al. (2020).
Temporal filtering also constitutes another relevant preprocessing step. From a general perspective, temporal filtering aims to remove portions of the frequency spectrum from the acquired fMRI data without relevant neurophysiological information (Poldrack et al., 2011). However, filtering is not without problems as it has been discussed in the literature (Lurie et al., 2020), even considering it inappropriate (Yuen et al., 2019). For example, several authors have pointed out that low-pass filtering can be problematic since it may remove signals of interest (Bolton et al., 2020a), while others have pointed out that high-pass filtering also has other downsides Lurie et al. (2020). In addition, the choice of filter parameters, such as cutoff frequencies, which can impact the analysis outcomes, is unclear since the optimal parameters vary among individuals. Consequently, finding an adequate filtering process that balances preserving relevant information and removing noise remains challenging in fMRI research.
@@
On the other hand, on top of the standard preprocessing steps, many approaches for FC analysis utilize some classical fMRI source separation techniques. One of the most relevant approaches is ICA, which plays a crucial role in many analysis pipelines. In addition to separating neuronal networks of interest, ICA is widely used for separating interfering components, e.g., cardiac or motion artifacts, from neurophysiologically relevant sources or as a data-driven way to spatially parcellate the brain into relevant brain sources. Formally, ICA is a statistical-based method that assumes that the fMRI data can be described as a linear combination of a set of maximally independent sources (Sergios Theodoridis, 2020). Additionally, although often overlooked (Lenton et al., 2023), ICA also assumes that the fMRI is stationary from a statistical point of view.
@@
Nevertheless, although there exist alternative approaches that rely on similar assumptions, there is compelling evidence of the brain’s nonlinear and nonstationary nature. Guan et al. (2020) explicitly studied the stationary and nonstationary behavior of different brain areas, showing that the brain dynamics are nonstationary and nonlinear. Furthermore, Zalesky et al. (2014) and Preti et al. (2017) reported consistent evidence supporting brain nonstationary behavior. As emphasized by Guan et al. (2020), researchers should consider nonlinear and nonstationary effects when studying brain FC.
Consequently, it becomes evident that there is a need for more advanced and adaptive preprocessing methods capable of incorporating the nonlinearity and nonstationarity nature of fMRI data. Considering all unique challenges brain activity poses, as we have discussed (Bolton et al., 2020b; Guan et al., 2020; Cordes et al., 2001), we propose a novel approach that separates neuronal brain activity from other interfering components and, at the same time, unravels those brain activity across different timescales, while naturally accounting for the dynamic, nonlinear, and multifaceted nature of brain activity, in a fully data-driven way tailored to each individual.
Put succinctly, our proposed methodology, referred to as Multiscale Functional Connectivity, relies on Multivariate Mode Decomposition (MMD) a Signal Processing technique that allows us to capture brain activity patterns at different timescales and separate them into neurophysiological components and interfering signals. To achieve this, MMD assumes that the fMRI data constitute a mixture of multiple inherent oscillatory components. Then, we use the results from this decomposition to study FC across different timescales.
In this regard, we would like to mention that the idea of decomposing fMRI data into their inherent oscillations has been explored by Yuen et al. (2019). However, unlike Yuen et al. (2019), who performed their study voxel-wise and utilized a band-pass filter for eliminating interfering components, we take some steps further by (a) incorporating the multivariate nature of fMRI, (b) avoiding any band-pass filter to separate the relevant components, and (c) exploring the FC associated with these oscillatory components further.
@@
In summary, we propose an alternative way for extracting FC information from fMRI data, bypassing the limitations of conventional preprocessing steps by computing the nonlinear, nonstationary, and multivariate nature of fMRI data. In addition to aligning with the expected natural complexities of the fMRI data, our proposed approach can detect and separate interfering components, such as structural noise from heart and respiration components, and simultaneously offers new insights into how the brain networks operate at different timescales. Finally, we evaluated the performance of our proposed methodology using three distinct fMRI studies. Our results demonstrate that our novel approach effectively separates the fMRI data into different timescales while identifying functional connectivity patterns with high reliability across individuals. Furthermore, we further extended our knowledge of how the FC for these three experiments spans among different timescales.
2 Materials and Methods
@@ 2.1 Natural frequencies in fMRI data
To fully appreciate the motivation behind the proposed study, we need a clear understanding of the frequency behavior of the fMRI data. Firstly, while studies focusing on frequency-related aspects of fMRI are relatively sparse, existing research offers valuable insight into the frequency organization of the fMRI signal and brain dynamics. For instance, Cordes et al. (2001) demonstrates that the frequency contribution to the correlation patterns spans several frequency bands. Similarly, Yuen et al. (2019) investigated the inherent frequency components across different brain locations –in a voxel-wise fashion– yielding similar findings.
Unlike other commonly used alternative neuroimaging techniques such as EEG or MEG, in fMRI, neuronal activity is indirectly measured through the Blood Oxygen Level Dependent (BOLD) effect (Power et al., 2011), which limits the observable neuronal activation dynamics of the brain to lower frequencies (Preti et al., 2017). Similarly, the relatively low sampling ratio utilized by fMRI scanners also imposes a challenge when studying the frequency contributions to fMRI signals (Power et al., 2011), as it restricts the maximum potential accessible frequencies and, at the same time, can introduce aliasing with physiological components (Cordes et al., 2001).
@@
Despite these limitations, fMRI frequency components comprise a rich spectrum that covers several relevant frequency bands. First, very low-frequency oscillations, lower than 10 mHz (Power et al., 2011), correspond to trends, scanner instabilities, and motion residuals. Neurophysiological activation patterns resulting from neuronal activity appear within the range of 10 to 200 mHz  (Cordes et al., 2001; Yuen et al., 2019) emphasizing the significant contribution of this frequency band to fluctuations related to brain activity, which corresponds with the natural band dominated by the BOLD response.

### v2

# Multiscale Functional Connectivity: Exploring the brain functional connectivity at different timescales

@@ Abstract
Human brains exhibit highly organized multiscale neurophysiological dynamics. Comprehending those dynamic changes and the neuronal networks involved is critical for understanding how the brain functions in health and disease. Functional Magnetic Resonance Imaging (fMRI) is a prevalent neuroimaging technique for studying these complex interactions. However, analyzing fMRI data poses several challenges. Most approaches for analyzing Functional Connectivity (FC) still rely on preprocessing or conventional methods, often built upon oversimplified assumptions. On top of that, those approaches often ignore frequency-related information despite evidence showing that fMRI data contain rich information that spans multiple timescales. This study introduces a novel methodology, Multiscale Functional Connectivity (MFC), to analyze fMRI data by decomposing the data into their intrinsic oscillatory component, allowing us to separate the neurophysiological activation patterns at multiple timescales while separating them from other interfering components. Additionally, the proposed approach accounts for the nonlinear, nonstationary, and multivariate nature of fMRI while allowing for the particularities of each individual in a data-driven way. We evaluated the performance of our proposed methodology using three fMRI experiments. Our results demonstrate that our novel approach effectively separates the fMRI data into different timescales while identifying highly reliable functional connectivity patterns across individuals.
{strip}
1 Introduction
@@
The brain is a complex system that exhibits highly organized neurophysiological interaction at various time scales [1, 2, 3]. As a result, brain activity displays highly complex spatiotemporal dynamic [3, 4], and their knowledge is crucial to enhancing our understanding of the brain’s functioning and cognition [5, 6, 7].
@@
Researchers have used a variety of neuroimaging techniques to study brain activity. Among all those techniques, functional Magnetic Resonance Imaging (fMRI) stands as one of the most popular for both neurophysiological, e.g., [8, 9, 10], and clinical research, e.g., [11]. Unlike other alternative noninvasive neuroimaging techniques, such as Electroencephalography (EEG), e.g., [12], or Magnetoencephalography (MEG) [13], fMRI records brain activity indirectly by measuring the variation in the oxygenation levels on small volumes of tissue, referred to as voxels [14], due to the metabolic oxygen consumption of the neurons, which is a process referred to as Bold Oxygenation Level-Dependent (BOLD) contrast [6]. In addition to the brain-induced activity, the fMRI data contains a mixture of several interfering signals, such as movement, respiratory, and cardiac pulsations [15].
During the last two decades, fMRI research has increasingly focused on exploring the brain’s functional connectivity (FC) and the dynamic interactions among neuronal networks, as envisioned by Friston back in 2011 [7]. This shift reflects a broader recognition of the brain’s complex organization beyond localized activations in response to stimuli or tasks [7, 16, 3]. These conventional approaches focus on unveiling the activation areas and their corresponding activation patterns in response to stimuli, often referred to as sources [5, 14, 17].
Meanwhile, in light of the success of Machine Learning and Artificial Intelligence in a wide range of complicated problems, current approaches for studying FC often leverage machine or deep learning-based models. These models have shown a great potential to uncover hidden interactions among brain networks that are not directly observable from the data [18]. Yet, these approaches still depend on some fundamental preprocessing steps or conventional fMRI analysis techniques, such as Independent Component Analysis (ICA) [16], which remain critical for their application in practice [19, 3]. These preliminary steps, however, present several limitations, which hinder the realization of the full potential of those more advanced approaches.
@@
For instance, conventional motion correction approaches rely on the parametric rigid-body model, which assumes that the brain only undergoes rigid motion. These approaches provide a reasonable estimation for main motion effects. Nonetheless, these motion components are richer than this simple model, as they can carry more information regarding certain behavioral and physiological aspects of the participants, such as arousal [20], and even age [21]. Furthermore, these simple models overlook other more subtle sources of motion, such as respiratory or cardiac pulsations. These additional motion components are often challenging to identify since they appear highly structured, individual dependent, and even mimic other brain sources [22].
Temporal static filtering constitutes another prevalent preprocessing step that aims to remove portions of the frequency spectrum from the acquired fMRI data with no relevant neurophysiological information [6, 4]. However, this step typically involves a predefined selection of filter parameters, such as the number of components, cut-off frequencies, and corresponding bandwidths, which are hard to estimate in practice, among other limitations, see [23, 3]. That said, using the signal’s frequency information to extract its neurophysiological components and remove unwanted interfering components holds promise and has already been investigated in multiple studies. For instance, [23] have recently explored the idea of decomposing the fMRI signal into their inherent oscillatory components. However, [23] performed their study voxel-wise, ignoring the multivariate nature of fMRI data, and they still used a band-pass filter to remove interfering components. Consequently, finding an adequate filtering process that balances preserving relevant information and removing noisy interfering components remains challenging in fMRI research.
@@
In addition to these preprocessing steps, many approaches for FC analysis still utilize some conventional fMRI source separation techniques as part of their preprocessing pipelines [4, 24]. One of the most relevant approaches is ICA [16], which plays a crucial role in many analysis pipelines by (a) separating neuronal networks of interest, (b) separating interfering components, e.g., cardiac or motion artifacts, from neurophysiologically relevant sources [25], and (c) as a data-driven way to parcellate the brain into relevant brain sources. Despite its wide use in fMRI data analysis, ICA assumes that the fMRI data is –from a statistical point of view– independent and stationary [^1] Formally, ICA is a statistical-based method that assumes that the data of interest can be described as a linear combination of a set of maximally independent sources. The assumption of statistical independence also implies that the data are weekly stationary [26]., an assumption often overlooked in practice and may not correspond with the natural brain behavior [24].
@@
Nevertheless, although alternative approaches rely on similar assumptions, there is compelling evidence of the brain’s nonlinear and nonstationary nature. Authors in [27] explicitly studied the stationary and nonstationary behavior of different brain areas, showing that the brain dynamics are nonstationary and nonlinear. Furthermore, [1] and [2] reported consistent evidence supporting brain nonstationary behavior. As emphasized in [27], researchers should consider nonlinear and nonstationary effects when studying brain FC.
Consequently, it becomes evident that there is a need for more advanced and adaptive preprocessing methods capable of incorporating the nonlinear and nonstationary nature of fMRI data [3, 27, 28]. By separating neuronal brain activity from other interfering components and unraveling neuronal activity across multiple frequency bands, we address these challenges. Our proposed approach, referred to as Multiscale Functional Connectivity (MFC), allows measuring the brain’s FC at different timescales without any static preprocessing or source separation, with the aid of the Multivariate Mode Decomposition (MMD) [29, 30, 31] model. In addition, it is fully data-driven, and it automatically tunes the parameters of the relevant frequency bands based on the input data.
@@
In summary, we propose an alternative way to extract FC information from fMRI data, bypassing some of the limitations of conventional preprocessing steps. In addition to aligning with the expected natural complexities of the fMRI data, our proposed approach can separate interfering components, such as structural noise from heart and respiration components, and simultaneously offers new insights into how the brain networks operate at different timescales. Finally, we evaluated the performance of our proposed methodology using three distinct fMRI studies. Our results demonstrate that our novel approach effectively separates the fMRI data into different timescales while identifying FC patterns with high reproducibility across individuals, and we further extended our knowledge of how the FC for these three experiments spans among different timescales.
2 Methods
2.1 Background: Multivariate Mode Decomposition
Formally, Multivariate Mode Decomposition (MMD) is a Signal Processing model that assumes that a multivariate signal of interest accepts a representation as a linear combination of a set of a particular family of amplitude- and frequency-modulated (AM-FM) functions, with a well-defined instantaneous frequency at any given time instance among all the channels [29, 31], which are formally defined as $c_{k}(t)=a_{k}(t)\cos(\phi_{k}(t))$, where $a_{k}(t)$ and $\phi_{k}(t)$ are the amplitude and instantaneous frequency functions associated with the $k$-th component. In this way, the MMD model decomposes a multivariate signal, say $\boldsymbol{x}(t)=[x_{1}(t),x_{2}(t),\ldots,x_{R}(t)]^{\mathsf{T}}$, where $R$ is the total number of channels, into a linear combination of intrinsic oscillatory components, referred to as Intrinsic Modes (IMs), as follows:
$\boldsymbol{x}(t)=\sum_{k=1}^{K}\boldsymbol{c}_{k}(t),$ (1)
where $\boldsymbol{c}_{k}$ is the $k$-th multivariate component and $K$ is the total number of components [31]. In other words, these IM functions behave similarly to harmonics that remain relatively close to a particular frequency. Yet they are flexible enough to accommodate fluctuations in both amplitude and frequency in a data-driven way [32].
Unlike other similar alternative studies, e.g., [23], MMD inherently exploits the multivariate nature of fMRI data. The multivariate nature of MMD allows each voxel or region of interest (ROI) within each specific IM to exhibit a distinct functional activation pattern while sharing a common oscillatory behavior, i.e., a common central frequency. This flexibility accommodates a wide range of response changes. This behavior also contrasts with conventional approaches, such as ICA, where each voxel/ROI activity is decomposed as a linear combination of a single common time activation pattern [16].
However, performing MMD constitutes a challenging task, and several methods have been proposed to solve it. Among all the available approaches, in this study, we will focus only on two of the most popular alternatives: Multivariate Empirical Mode Decomposition (MEMD) [30] and Multivariate Variational Mode Decomposition (MVMD) [31]. Nevertheless, although both algorithms aim to achieve the same signal decomposition in a fully data-driven fashion, they are vastly different from an algorithmic perspective.
MEMD.
MEMD obtains intrinsic modes, $c_{k}(t)$ in (1), using an iterative sifting process. The sifting process unfolds by iteratively estimating the local mean of a signal by averaging its maximum and minimum envelope signals. The local mean signal is, by design, a low-frequency signal that is subtracted from the original signal to obtain a candidate for an IM. The drawback of MEMD is that it is highly sensitive to noise and suffers from the mode-mixing problem –a single mode carrying fundamentally different frequencies [33].
MVMD.
MVMD [^2] The implemented code for the MVMD algorithm is openly available on the MATLAB file exchange: https://se.mathworks.com/matlabcentral/fileexchange/72814-multivariate-variational-mode-decomposition-mvmd obtains a similar decomposition as given in (1) but unlike MEMD, uses an optimization formulation to estimate all intrinsic modes of a signal simultaneously. The objective is to find an estimate of narrowband functions from the data, where the center frequency of each mode aligns across all the multivariate components of the signal. This renders MVMD more robust to noise and provides a better estimate of the modes [31].
@@ 2.2 Natural frequencies in fMRI data
Understanding the frequency organization of the fMRI signal and brain activity dynamics is crucial for MFC. While studies focusing on frequency-related aspects of fMRI are relatively sparse, existing research offers valuable insight into the frequency organization of the fMRI signal and brain dynamics. For instance, [28] demonstrates that the frequency contribution to the correlation patterns spans several frequency bands. Similarly, [23] investigated the inherent frequency components across different brain locations –in a voxel-wise fashion– yielding similar findings.
@@
Overall, fMRI frequency components comprise a rich spectrum that covers several relevant frequency bands. First, very low-frequency oscillations, lower than 10 mHz [14], correspond to trends, scanner instabilities, and motion residuals. Neurophysiological activation patterns resulting from neuronal activity appear within the range of 10 to 200 mHz  [28, 23, 4] emphasizing the significant contribution of this frequency band to fluctuations related to brain activity, which corresponds with the natural band dominated by the BOLD response.